# A Universal Bound on QCD Axions from Supernovae

### v1

@@ I Introduction
In recent years, there has been increased interest in the physics of the QCD axion [1, 2, 3, 4], supported by ongoing experimental efforts and new theoretical ideas. While experiments increased their sensitivity enormously over the last decade, the strongest bounds still come from astrophysical probes, among which neutron star (NS) and supernova (SN) cooling are the most sensitive [5, 6].
@@
The first term in Eq. (3) is negligible for particle production since it results in an axion-nucleon coupling involving at least two axions, which is thus suppressed by $1/f_{a}^{2}$.  [^1] Note that this term has been extensively used in the literature as it can lead to the sourcing of light scalar fields in dense and large objects [54, 62, 63, 64, 65, 58, 66, 60].
@@ II Axion Production Channels
where $m_{N}$ denotes the nucleon mass, $F^{\mu\nu}$ is the photon field strength tensor, and $|C_{aN\gamma}|\simeq 10^{-3}$ as determined in the seminal work of [19], see also [20].
@@
The next-to-leading order correction to Fig. 2 comes from diagrams shown in Fig. 3, where the axion is attached to an external nucleon line by the operator
@@
where $S^{\mu}=\frac{i}{2}\gamma_{5}\sigma^{\mu\nu}v_{\nu}$ is the nucleon spin operator with $\sigma^{\mu\nu}=\frac{i}{2}\left[\gamma^{\mu},\gamma^{\nu}\right]$ and $v_{\mu}=(1,\vec{0})^{T}$  [^3] Note that we are slightly deviating from the basis choice of e.g. [67] in that we split $\chi_{-}$ into its trace and its traceless part. . The axion appears in the trace contribution, and expanding to leading order, one finds
@@
In analogy to [29], we assume that the undetermined low-energy constant $\tilde{d}_{19}$ is drawn from a superposition of two normal distributions with $\tilde{d}_{19}=\pm 0.5(5)~{}\text{GeV}^{-2}$. This is estimated using naive-dimensional-analysis and matches the measured magnitude of the traceless version of the operator in Eq. (7). The corrections coming from these diagrams are suppressed by one power of $k/\Lambda$ compared to the leading order contribution. We include all correction to this order, and estimate the effect of the truncation of the ChPT expansion for next-to-next-to leading order in analogy to [10].
Figure 3: Tree-level NLO axion production diagrams from the Lagrangian Eq. (7). The empty circle is the NLO axion-nucleon vertex. The LHS show the one-pion exchange while the RHS is the leading order contact diagram proportional to $C_{S}$, for details see [10]. These diagram is representative of the 8 diagrams that arise from permutations of the axion vertex connected to a nucleon line $N=(p,n)^{\rm T}$.
In the following we show the leading order squared matrix elements resulting Fig. 2 for the two cases of interest: identical nucleons in the initial state ($nn$ or $pp$) and the mixed case ($np$). For identical external nucleons, there are eight diagrams consisting of the two possible combinations of external states, i.e. , $pp\to ppa$ and $nn\to nna$, each with $t-$ and $u-$channel, and two possibilities of attaching the axion vertex to the diagram. They give rise to the squared matrix element
$\left\lvert\mathcal{M}_{nn/pp}\right\rvert^{2}=\frac{4608\,\hat{c}_{5}^{2}\,g_{A}^{2}\,m_{N}^{4}\,m_{\pi}^{4}}{f_{a}^{2}\,f_{\pi}^{4}}\frac{z^{2}}{(z+1)^{4}}\,\,\times$ (9)
$\Bigg{[}\frac{\mathbf{k}^{2}}{\left(\mathbf{k}^{2}+m_{\pi}^{2}\right)^{2}}+\frac{\mathbf{l}^{2}}{\left(\mathbf{l}^{2}+m_{\pi}^{2}\right)^{2}}-\frac{2\left(\mathbf{k}\cdot\mathbf{l}\right)}{\left(\mathbf{k}^{2}+m_{\pi}^{2}\right)\left(\mathbf{l}^{2}+m_{\pi}^{2}\right)}\Bigg{]}.$
@@ III Axion Luminosity and Bound
×$\quad\times S|\mathcal{M}|^{2}E_{a}f_{1}f_{2}\left(1-f_{3}\right)\left(1-f_{4}\right).$
@@
[…]
@@
¯$\bar{\ell}_{3}$ [37] 3.41(82)
¯$\bar{\ell}_{4}$ [37] 4.40(28)
$\hat{c}_{1}$ [38] $-1.07(2)~{}\mathrm{GeV}^{-1}$
$\hat{c}_{5}$ [39] $-0.09(1)~{}\mathrm{GeV}^{-1}$
$C_{S}$ [40] $-119(7)~{}\mathrm{GeV}^{-2}$
@@ IV Conclusions
V Acknowledgments
We would like to thank Thomas Janka and his group for providing us with supernova simulation data. This work has been supported by the Collaborative Research Center SFB1258, the Munich Institute for Astro-, Particle and BioPhysics (MIAPbP), and by the Excellence Cluster ORIGINS, which is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC 2094 – 39078331. The research of MS is partially supported by the International Max Planck Research School (IMPRS) on “Elementary Particle Physics”. KS is supported by a research grant from Mr. and Mrs. George Zbeda. SS is partially supported by the Swiss National Science Foundation under contract 200020-213104. SS thanks the CERN theory group for its hospitality.

### v2

@@ I Introduction
In recent years, there has been increased interest in the physics of the QCD axion [1, 2, 3, 4], supported by ongoing experimental efforts and new theoretical ideas. While experiments increased their sensitivity enormously over the last decade, the strongest bounds still come from astrophysical probes, among which neutron star (NS) and supernova (SN) cooling are the most sensitive [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25].
@@
The first term in Eq. (3) is negligible for particle production since it results in an axion-nucleon coupling involving at least two axions, which is thus suppressed by $f_{a}^{2}. 11$1Note that this term has been extensively used in the literature as it can lead to the sourcing of light scalar fields in dense and large objects [66, 74, 75, 76, 68, 70, 77, 72].
@@ II Axion Production Channels
where $\sigma^{\mu\nu}=\frac{i}{2}\left[\gamma^{\mu},\gamma^{\nu}\right]$, $m_{N}$ denotes the nucleon mass, $F^{\mu\nu}$ is the photon field strength tensor, and $|C_{aN\gamma}|\simeq 10^{-3}$ as determined in the seminal work of [32], see also [33].
@@
The next-to-leading order correction to Fig. 2 comes from diagrams shown in Fig. 3, where the axion is attached to an external nucleon line by the model-independent operator
@@
where $S^{\mu}=\frac{i}{2}\gamma_{5}\sigma^{\mu\nu}v_{\nu}$ is the nucleon spin operator and $v_{\mu}=(1,\vec{0})^{T}$  [^3] Note that we are slightly deviating from the basis choice of e.g. [78] in that we split $\chi_{-}$ into its trace and its traceless part. . The axion appears in the trace contribution, and expanding to leading order, one finds
@@
In analogy to [42], we assume that the undetermined low-energy constant $\tilde{d}_{19}$ is drawn from a superposition of two normal distributions with $\tilde{d}_{19}=\pm 0.5(5)~\text{GeV}^{-2}$. This is estimated using naive-dimensional-analysis and matches the measured magnitude of the traceless version of the operator in Eq. (7). The corrections coming from these diagrams are suppressed by one power of $k/\Lambda$ compared to the leading order contribution. We include all corrections to this order, and estimate the effect of the truncation of the ChPT expansion for next-to-next-to leading order in analogy to [29]. We neglect phenomenological effects taken into account in [18, 19, 28] in order not to spoil the consistent systematics of our calculation.
Figure 3: Tree-level NLO axion production diagrams from the Lagrangian Eq. (7). The empty circle is the NLO axion-nucleon vertex. The LHS show the one-pion exchange while the RHS is the leading order contact diagram proportional to $C_{S}$, for details see [29]. This diagram is representative of the 8 diagrams that arise from permutations of the axion vertex connected to a nucleon line $N=(p,n)^{\rm T}$.
In the following, we show the leading order squared matrix elements resulting from Fig. 2 for the two cases of interest: identical nucleons in the initial state ($nn$ or $pp$) and the mixed case ($np$). For identical external nucleons, there are eight diagrams consisting of the two possible combinations of external states, i.e. , $pp\to ppa$ and $nn\to nna$, each with $t$- and $u$-channel, and two possibilities of attaching the axion vertex to the diagram. They give rise to the squared matrix element
$\hskip-8.53581pt\left\lvert\mathcal{M}_{nn/pp}\right\rvert^{2}=\frac{4608\,\hat{c}_{5}^{2}\,g_{A}^{2}\,m_{N}^{4}\,m_{\pi}^{4}}{f_{a}^{2}\,f_{\pi}^{4}}\frac{z^{2}}{(z+1)^{4}}\,\,\times$ (9)
$\hskip-8.53581pt\Bigg{[}\frac{\mathbf{k}^{2}}{\left(\mathbf{k}^{2}+m_{\pi}^{2}\right)^{2}}+\frac{\mathbf{l}^{2}}{\left(\mathbf{l}^{2}+m_{\pi}^{2}\right)^{2}}-\frac{2\left(\mathbf{k}\cdot\mathbf{l}\right)}{\left(\mathbf{k}^{2}+m_{\pi}^{2}\right)\left(\mathbf{l}^{2}+m_{\pi}^{2}\right)}\Bigg{]}.$
@@ III Axion Luminosity and Bound
$\quad\times S|\mathcal{M}|^{2}E_{a}f_{1}f_{2}\left(1-f_{3}\right)\left(1-f_{4}\right).$
@@
$L_{a}=\int dr4\pi r^{2}\dot{\varepsilon}_{a}(r)\text{(lapse)}^{2}(1+2v_{r}).$ (13)
We account for gravitational redshift and corrections from integrating the emissivity in the comoving frame, resulting in an additional radial-dependent factor of $\text{(lapse)}^{2}(1+2v_{r})$, as described in [45, 46, 43].
@@
$\bar{\ell}_{3}$ [49] 3.41(82)
$\bar{\ell}_{4}$ [49] 4.40(28)
$\hat{c}_{1}$ [50] $1.07(2)~\mathrm{GeV}^{-1}$
$\hat{c}_{5}$ [51] $0.09(1)~\mathrm{GeV}^{-1}$
$C_{S}$ [52] $119(7)~\mathrm{GeV}^{-2}$
@@ IV Conclusions
Acknowledgements.
We would like to thank Thomas Janka and his group for providing us with supernova simulation data. This work has been supported by the Collaborative Research Center SFB1258, the Munich Institute for Astro-, Particle and BioPhysics (MIAPbP), and by the Excellence Cluster ORIGINS, which is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC 2094 – 390783311. The research of MS is partially supported by the International Max Planck Research School (IMPRS) on “Elementary Particle Physics”. KS is supported by a research grant from Mr. and Mrs. George Zbeda. SS is partially supported by the Swiss National Science Foundation under contract 200020-213104. SS thanks the CERN theory group for its hospitality.
@@ Appendix A Alternative process
Appendix B Dependence on the Supernova Model
While the model-independent supernova bound put forward in this paper is independent of the specific QCD axion model, it does depend on the specifics of the supernova model. In this appendix, we study some aspects of the supernova model dependence on the bound derived.
Figure 8: Density and temperature profile taken from a supernova simulation of [43] at $t_{pb}=1\,\mathrm{s}$ for the three supernova models discussed in this section.
In the main text, we have studied the axion luminosity resulting from one specific supernova model based on the SFHo EOS that included muons with a progenitor star of a mass of $18.6\,M_{\odot}$ from [43]. In this appendix, we extend the analysis by two different supernova models, based on the same EOS but different progenitor masses and simulations, and perform the same calculation, again at LO. The SN simulations we take additionally into account have a $18.8\,M_{\odot}$ and $20\,M_{\odot}$ progenitor masses, both with the SFHo EOS and including muons, see [43]. While the quantitative numerical results differ for each of the models by an order of one factor, qualitatively, the results are in agreement.
The supernova profiles resulting from a $18.8\,M_{\odot}$ and a $20\,M_{\odot}$ progenitor from different simulations lead to a lower and a higher temperature profile, respectively. The resulting temperature profiles for the three simulations as a function of the radius at $t_{pb}=1\,\mathrm{s}$ are shown in Fig. 8. Calculating the resulting axion luminosity as presented in the main text at LO, we find the results shown in Fig. 9.
Figure 9: Axion luminosity $L_{a}f_{a}^{2}$ as a function of time for different contributions as shown in Fig. 4. The dashed lines show the LO emissivities, resulting from the diagram shown in Fig. 2, for different supernova models.
While the results for the individual supernovae are quite different, especially at late times, we still find for all of them a strengthening of the supernova bound by around two orders of magnitude compared to [31]. The axion luminosity from the interaction taken into account in [31] is only shown for the $18.6\,M_{\odot}$ progenitor mass. Since the relative effect from changing the SN model is identical to the change in luminosity with our interaction (blue, green dashed lines compared to black dashed line), they are suppressed in Fig. 9. The supernova with slightly higher temperatures, where the progenitor has a mass of $20\,M_{\odot}$, leads to a strengthening of the LO bound by a factor of around $2.7$, while the supernova with slightly lower temperatures and a progenitor of a mass of $18.8\,M_{\odot}$, leads to a weakening of the LO bound by a factor of around $0.6$. Including also the NLO contributions and performing the same statistical analysis as above, similar results are to be expected.